3.Consider the following information to answer questions

of 10% and to issue new preferred stock with a $4.00 pershare dvidend at $25 a share. The common

X Company is evaluating its cost of capita under alternative financing arrangements. In consultation

1 20% 30% 50%

2 50% 30% 20%

$2.50 per share next year. Market anaysts foresee a growth in dividends in Invest stock at a rate of

5% per year.nvest does not expect its cost of debt, preferred stock or common stock, to be different

Hint:coupon rate of the bond is the same as the beforetax cost of debt.

Percentage of New Capital

with nvestment bankers, X Company expects to be abe to issue new debt at par with a coupon rate

arrangement?

A.What s the weighted average cost of capital to X Company under the first financing arrangement?

stock of X Company is currenty seling for $20.00 a share. X Company expects to pay a dividend of

under the two possible financing arrangements. X Company margina tax rate is 40%.



1
Expert's answer
2022-01-04T10:20:52-0500

Let the bond rate be 10%. the share of preferred shares is 20%, ordinary shares 50%, debt 30%:

"WACC=0.2\times\frac{4}{25}\times100+0.5(\frac{2.5}{20}\times100+5)+0.3\times10(1-0.4)=13.75"